# Algebraic non-Hermitian skin effect and unified non-Bloch band theory in arbitrary dimensions

### v1

@@ II.1 Hermitian systems
For a specific eigenenergy $E$, the dispersion relation $\epsilon_{n}(k_{1},\ldots,k_{d})=E$ demarcates a $(d-1)$-dimensional subspace within the $d$-dimensional BZ, namely the equal-energy manifold. In this manuscript, we adopt the terminology from Fermi liquid theory and refer to this equal-energy manifold as the Fermi surface, which inherently possesses a dimensionality $\dim{\text{FS}}=\dim{\text{BZ}}-\dim{E}=d-1$. As will be shown below, this concept of Fermi surface plays an important role in the construction of eigenstates of $H$. This is because for each given eigenenergy, the corresponding eigenstate can be written as the superposition of Bloch waves in the corresponding Fermi surface
@@
Although embedded in a 2D momentum space, the actual dimension of GBZ is in fact less than two. This reduction results from the boundary conditions imposing extra constraints on the permitted $\beta$ values, thereby diminishing the GBZ’s dimensionality. The construction of the GBZ for 1D systems is already well-established, so we only provide a succinct example here. Consider a standard 1D tight-binding model, in which the Bloch Hamiltonian $\mathcal{H}(k)$ is analytically continued into its non-Bloch counterpart with the substitution $e^{ik}\to\beta$:
@@
where $t_{i}$ denotes generally complex coefficients, with $t_{0},t_{M+N}\neq 0$. The non-Bloch Hamiltonian $\mathcal{H}(\beta)$ thus becomes a Laurent polynomial in the complex variable $\beta$. For a given complex eigenenergy $E$, by solving the characteristic equation $\det[\mathcal{H}(\beta)-E]=0$, typically order-1 non-Bloch wave solutions $\beta_{i}(E)$ will be obtained, which can be arranged by their magnitudes as $|\beta_{i}(E)|\leq|\beta_{i+1}(E)|$. Apart from the characteristic equation, boundary conditions at the two ends of the 1D chain impose another important constraint, known as the standing wave condition
@@
which stands as a cornerstone of 1D GBZ theory Yao and Wang (2018); Yokomizo and Murakami (2019); Zhang et al. (2020); Yang et al. (2020); Kawabata et al. (2020a). This condition ensures the formation of standing waves, which can simultaneously satisfy the OBCs at both ends, and thus allows the emergence of an extensive number of edge states, each of which is composed of two plane waves of exponential form with identical localization length, i.e., the non-Hermitian skin effect. Since the standing wave condition introduces one restriction on the values of $\beta$, the space encompassed by the allowable $\beta$ values experiences a one-dimensional reduction, and thus $\dim{\text{GBZ}}=2d-1=1$.
Once the 1D GBZ is established, for each point within it (i.e., each permissible value of $\beta$), the corresponding eigenenergies $E$ can be inferred from the characteristic equation $\det[\mathcal{H}(\beta)-E]=0$. As the allowable $\beta$ values constitute a one-dimensional continuum, it follows that the admissible $E$ values also create a one-dimensional continuum, corresponding to a set of 1D segments in the complex energy plane, that is, $\dim{E}=1$.
@@
In 1D, the dimensionality of the GFS is determined by $\dim{\text{GFS}}=\dim{\text{GBZ}}-\dim{E}=0$. This zero dimensionality implies that the GFS consists of only a discrete set of $\beta$. This set can be calculated by solving the characteristic equation while satisfying the standing wave condition in Eq. (10) simultaneously. In such cases, the wavefunction $\Psi_{E}$ becomes the sum of a finite number of exponential functions in terms of $x$. Consequently, $\Psi_{E}$ must decay exponentially at large $x$, if it is not a bulk Bloch wave. The fact that $\dim\text{GFS}=0$ in 1D is the reason why all skin modes decay exponentially in 1D non-Hermitian skin effects.
@@ II.3 Non-Hermitian Systems in $d>1$
1 number of GBZ $1\leq\text{ number of GBZ constraints}\leq d.$ (15)
[…]
@@
Regarding the OBC eigenenergies, non-Hermitian systems admit complex energy eigenvalues. Plotting all permissible energy eigenvalues on the complex energy plane, two scenarios may emerge. In the most generic scenario, these OBC eigenenergies will occupy certain 2D areas of the complex energy plane, i.e., the energy spectrum is two-dimensional ($\dim{E}=2$). This represents a marked departure from Hermitian systems or 1D non-Hermitian systems, where $\dim{E}=1$. In an alternative scenario, which may arise from symmetry, fine-tuning, or other factors, the allowed OBC energy values fail to span any 2D regions on the complex energy plane. Instead, they are confined to 1D curves, and hence $\dim{E}=1$.
@@
where the summation of $\beta$ extends over all possible values within the generalized Fermi surface of $E$. The dimension of the generalized Fermi surface is determined by the relation $\dim{\text{GFS}}=\dim{\text{GBZ}}-\dim{E}$. Therefore, the dimension of the generalized Fermi surface may vary from 00 up to $22d-$ 2. Unlike 1D non-Hermitian systems, which are characterized by a 0-dimensional generalized Fermi surface, systems with $d>1$ may exhibit $\dim{\text{GFS}}>0$. The presence of a nonzero $\dim{\text{GFS}}$ indicates that the summation in Eq. (17) becomes an integration under the thermodynamic limit with OBCs. Because integrals of exponential functions can engender nearly any function form (for instance, through the Laplace transform), even though the eigenstate is built from exponential functions of the coordinates $(x_{1},\ldots,x_{d})$ [Eq. (17)], it can manifest extremely diverse behaviors in non-Hermitian systems where $d>1$. For example, the eigenstate may turn into a power-law function of the coordinates, as demonstrated in Sec. IV. It is this richness that underpins the potential for the algebraic skin effect exclusive to non-Hermitian systems in dimensions higher than one.
@@
II.4 Strategy to formulate the universal GBZ formula
Here, we succinctly outline our approach for deriving the GBZ formula for non-Hermitian systems in arbitrary dimensions, with comprehensive formalism to be detailed in the subsequent sections.
Consider a system of dimension $d$, with the system size $L_{1}\times L_{2}\times\ldots\times L_{d}$. We can label the unit cells using integer coordinates $(x_{1},x_{2},\ldots,x_{3})$ with $1\leq x_{i}\leq L_{i}$. The transfer matrix method is employed to construct the eigenstates of the system — an exact analytic technique that systematically yields all eigenenergies and eigenstates for a given Hamiltonian Lee and Joannopoulos (1981); Hatsugai (1993); Dwivedi and Chua (2016); Kunst and Dwivedi (2019). This method has been widely used for Hermitian systems under both open and periodic boundary conditions, and it is equally effective for non-Hermitian systems.
@@ II.4 Strategy to formulate the universal GBZ formula
Here the bulk Hamiltonian is encoded in $\mathbb{T}$. For a fixed $\beta_{d}$, the reduced subsystem can be treated as a $(d-1)$-dimensional system, $\psi_{\beta_{d}}(x_{1},\ldots x_{d-1})$, and thus we can solve this $(d-1)$-dimensional problem using the transfer matrix method once again, defining a transfer matrix for $x_{d-1}$:
@@
Iterating this process ultimately simplifies the issue to a one-dimensional setup along the direction of $x_{1}$, $\psi_{\beta_{d}\beta_{d-1}\dots\beta_{2}}(x_{1})$. In this way, the $d$-dimensional system is reduced into a 1D system of $\psi_{\beta_{d}\beta_{d-1}\dots\beta_{2}}(x_{1})$, supplemented by transfer matrices for the other $d-1$ directions which only rely on bulk properties. We can then treat this 1D problem as a 1D non-Hermitian system, where the GBZ theory is well understood: the non-Hermitian skin effect necessitates a standing wave condition for the direction of $x_{1}$. This condition ensures that standing waves can be formed to obey boundary conditions at $x_{1}=1$ and $x_{1}=L_{1}$, allowing extensive number of skin modes. However, once this condition is enforced along $x_{1}$, imposing additional standing-wave constraints in other directions is generally untenable, which will be discussed below. Thus, for other directions, we must allow all possible eigenvalues of the transfer matrices $\mathbb{T}_{2},\ldots,\mathbb{T}_{d}$ without further stipulations.
@@
Since the standing wave condition can typically be imposed in just one spatial direction, as previously discussed, the GBZ’s complex wavevector space $(\beta_{1},\ldots,\beta_{d})$ characteristically spans a space of dimension $12d-$ 1, surpassing real space dimensions $d$ if $d>1$. This distinctive characteristic underpins why non-Hermitian systems with dimensions greater than 1D can exhibit unprecedented phenomena.
@@
Figure 1:  Dimension analysis for GBZ and GFS in 1D and 2D systems. The GBZ dimension in $d$-dimensional systems ranges from $d$ to $12d-$ 1. In 1D systems, $\dim{\text{GBZ}}=1$, and the OBC spectrum (denoted by $\sigma_{\text{OBC}}$) is limited to forming arcs, meaning $\dim{\sigma_{\text{OBC}}}=1$. This leads to $\dim{\text{GFS}}=0$ based on Eq. (5), with OBC eigenstates either being bulk or exponentially localized wavefunctions. In 2D systems, $\dim{\text{GBZ}}$ can be $2$ or $3$, corresponding to the standing wave conditions that are enforced in two or one spatial dimensions, respectively. The OBC spectrum in 2D generally spans a finite area in the complex energy plane, $\dim{\sigma_{\text{OBC}}}=2$. Consequently, $\dim{\text{GBZ}}=0$ or $1$. When $\dim{\text{GFS}}=0$, the OBC eigenstates behave as bulk or exponentially localized wavefunctions. When $\dim{\text{GFS}}>0$, the OBC eigenstates can exhibit power-law decaying behavior, a phenomenon unique to high dimensions.
@@
As we conclude this section, it is important to draw attention to two key points. The first point is regarding the application of the standing wave condition, which can be imposed in an arbitrary direction. This flexibility implies that, depending on the choice of standing-wave-condition directions, $d$ distinct GBZs can be defined, and each GBZ corresponds to a unique set of basis functions. Because standing wave conditions in different directions in general produces different basis functions, they are typically not compatible with each other and thus cannot be simultaneously enforced. However, it is essential to recognize that while different bases associated with different GBZs can be used, the resulting eigenvalue spectrum and eigenstates—constructed as superpositions of these basis functions—remain invariant with respect to the choice of GBZ. Put simply, although multiple GBZ bases can be defined, they all ultimately describe the same physical reality. Secondly, although typically only one direction can accommodate the standing wave condition, special scenarios may allow the standing-wave conditions across multiple spatial directions, e.g., decoupled 1D chains populating a $d$-dimensional space. In such systems, GBZ dimension is further reduced to below $12d-$ 1. If standing wave conditions for all directions are compatible with each other, the dimension of the GBZ reaches its minimum possible value, which coincides with the real-space dimensions $d$. This framework for 1D and 2D systems is encapsulated in Fig. 1.
@@
III Non-Bloch band theory in two dimensions based on the layer transfer matrix method
@@
In this section, we offer a general recipe for deriving the universal GBZ formula of 2D lattice systems using a layer transfer matrix approach. The extension of the GBZ formula to higher dimensions ($d>2$) is detailed in Section V.
@@
To construct OBC eigenstates, we begin with the layer transfer matrix for a given tight-binding Hamiltonian. The eigenvalues of the transfer matrix dictate the propagation of the wavefunction’s layer components along the transfer matrix direction. We demonstrate that a generic 2D OBC eigenstate can be represented by a superposition of standing wave components along one spatial direction (e.g., the $x$ direction), each coupled with the corresponding allowable complex momentum $\beta_{y}$ in the $y$ direction. It is expressed as: $\psi(x,y)=\sum_{i}A_{i}\,\rho^{y}_{y,i}\phi_{i}(x)$, where $\phi_{i}(x)$ represents $i$-th standing wave along the $x$ direction and $\rho_{y,i}$ is the transfer matrix eigenvalue associated with $\phi_{i}(x)$. A typical standing-wave component of an OBC eigenstate is illustrated in Fig. 2, where two equal-amplitude and counter-propagating complex waves $\beta_{x,i}$ and $\widetilde{\beta}_{x,i}$ form standing wave $\phi_{i}(x)$ in the $x$ direction and transfer along the $y$ direction characterized by $\rho_{y,i}$.
From this construction recipe for OBC eigenstates, the formation of standing waves in either one or multiple directions necessitates the corresponding GBZ conditions [Eq. (14)]. We can obtain the GBZ manifold according to the GBZ conditions. As a verification of our GBZ formula, we construct the OBC wavefunctions from the GBZ basis, which perfectly agrees with the numerical simulations.
Since the layer transfer matrix depends on the choice of layer basis, for a generic 2D Hamiltonian, the GBZ condition (or standing-wave condition) and the resultant GBZ manifold are not unique. This is fundamentally distinct from the 1D cases. However, it is worth noting that the distinct GBZ manifold only implies a basis gauge redundancy: although the GBZ manifold is not unique and depends on the layer basis, the same physical consequences and quantities will ultimately be derived.
We will detail the above main findings and conclusions in this section organized as follows: Subsections III.1III.3 focus on the GBZ formalism for the most generic cases in 2D, where $\dim{\sigma_{\text{OBC}}}=2$ and $\dim{\text{GBZ}}=3$, as illustrated in Fig. 1. We then remark on the non-uniqueness of higher dimensional GBZ manifold and GBZ conditions in subsection III.4. Finally, we discuss the GBZ formula for other 2D scenarios in subsection III.5.
@@
Figure 2:  Schematic of $i$-th standing-wave component of an OBC eigenstate under the layer-$y$ transfer matrix basis. Here, the $i$-th component is a standing wave $\phi_{i}(x)$ in the $x$ direction, formed by two equal amplitude and counter-propagating non-Bloch waves $\beta_{x,i}$ and $\widetilde{\beta}_{x,i}$, coupled with the associated $\rho_{y,i}$ in the $y$ direction.
III.1 Bulk: the non-Bloch layer transfer matrix

### v2

@@ II.1 Hermitian systems
For a specific eigenenergy $E$, the dispersion relation $\epsilon_{n}(k_{1},\ldots,k_{d})=E$ demarcates a (d1)(d-1)-dimensional subspace within the $d$-dimensional BZ, namely the equal-energy manifold. In this paper, we adopt the terminology from Fermi liquid theory and refer to this equal-energy manifold as the Fermi surface, which inherently possesses a dimensionality $\dim{\text{FS}}=\dim{\text{BZ}}-\dim{E}=d-1$. As will be shown below, this concept of Fermi surface plays an important role in the construction of eigenstates of $H$. This is because for each given eigenenergy, the corresponding eigenstate can be written as the superposition of Bloch waves in the corresponding Fermi surface
@@
Although embedded in a 2D momentum space, the actual dimension of GBZ is in fact less than two. This reduction results from the open boundary conditions, where the standing-wave condition imposes extra constraints on the permitted $\beta$ values, thereby diminishing the GBZ’s dimensionality. The construction of the GBZ for 1D systems is already well-established, so we only provide a succinct example here. For simplicity, consider a standard 1D single-band tight-binding model, in which the Bloch Hamiltonian $\mathcal{H}(k)$ is analytically continued into its non-Bloch counterpart with the substitution $e^{ik}\to\beta$:
@@
where $t_{i}$ denotes generally complex coefficients, with $t_{0},t_{M+N}\neq 0$. The non-Bloch Hamiltonian $\mathcal{H}(\beta)$ thus becomes a Laurent polynomial in the complex variable $\beta$. For a given complex eigenenergy $E$, by solving the characteristic equation $\det[\mathcal{H}(\beta)-E]=0$, typically order-1 non-Bloch wave solutions $\beta_{i}(E)$ will be obtained, which can be arranged by their magnitudes as $|\beta_{i}(E)|\leq|\beta_{i+1}(E)|$. Apart from the characteristic equation, boundary conditions at the two ends of the 1D open chain impose another important constraint, known as the standing wave condition
@@
which stands as a cornerstone of 1D GBZ theory [25, 27, 30, 32, 33]. For further conclusions on generic multi-band models, see Refs.[27, 33]. This condition ensures the formation of standing waves, which can simultaneously satisfy the OBCs at both ends, and thus allows the emergence of an extensive number of edge states, each of which is composed of two plane waves of exponential form with identical localization length, i.e., the non-Hermitian skin effect. Since the standing wave condition introduces one restriction on the values of $\beta$, the space encompassed by the allowable $\beta$ values experiences a one-dimensional reduction, and thus $\dim{\text{GBZ}}=2d-1=1$.
Once the 1D GBZ is established, for each point within it (i.e., each permissible value of $\beta$), the corresponding eigenenergies $E$ can be inferred from the characteristic equation $\det[\mathcal{H}(\beta)-E]=0$. As the allowable $\beta$ values constitute a one-dimensional continuum, it follows that the admissible $E$ values also create a one-dimensional continuum, corresponding to a set of 1D segments in the complex energy plane, that is, $\dim{E}=1$, as illustrated in Fig. 2.
@@
In 1D, the dimensionality of the GFS is determined by $\dim{\text{GFS}}=\dim{\text{GBZ}}-\dim{E}=0$. This zero dimensionality implies that the GFS consists of only a discrete set of $\beta$. This set can be calculated by solving the characteristic equation while satisfying the standing wave condition in Eq. (10) simultaneously. In such cases, the wavefunction $\Psi_{E}$ becomes the sum of a finite number of exponential functions in terms of $x$. Consequently, $\Psi_{E}$ must decay exponentially at large $x$, if it is not an extended Bloch wave. The existing GBZ framework implies that for a given OBC eigenvalue, the localization length is uniquely determined, resulting in the exponential localization characteristic of the 1D non-Hermitian skin effect.
@@ II.3 Non-Hermitian Systems in $d>1$
$1\leq\text{ number of GBZ constraints}\leq d.$ (15)
The lower limit, $\text{number of GBZ constraints}=1$, indicates that standing wave conditions for any two distinct spatial directions are not reconcilable, thereby allowing for just one such condition for a single spatial direction. The upper limit, $\text{number of GBZ constraints}=d$, suggests that standing wave conditions for all distinct directions are congruent, permitting their simultaneous enforcement. Drawing from Eq. (13), at the lower limit of standing wave conditions ($\text{number of GBZ constraints}=1$), the GBZ dimension reaches 2d12d-1, while at the upper limit ($\text{number of GBZ constraints}=d$), the GBZ dimension is reduced to $d$. Typically, a non-Hermitian system will fall between these extremes, hence the dimension of the GBZ lies between $d$ and 2d12d-1:
@@
Regarding the OBC eigenenergies, non-Hermitian systems admit complex energy eigenvalues. Plotting all permissible energy eigenvalues on the complex energy plane, two scenarios may emerge. In the most generic scenario, these OBC eigenenergies will occupy certain 2D areas of the complex energy plane, i.e., the energy spectrum is two-dimensional ($\dim{E}=2$), as shown in Fig. 2. This represents a marked departure from Hermitian systems or 1D non-Hermitian systems, where $\dim{E}=1$. In an alternative scenario, which may arise from symmetry, fine-tuning, or other factors, the allowed OBC energy values fail to span any 2D regions on the complex energy plane. Instead, they are confined to 1D curves, and hence $\dim{E}=1$.
@@
where the summation of $\beta$ extends over all possible values within the generalized Fermi surface of $E$. The dimension of the generalized Fermi surface is determined by the relation $\dim{\text{GFS}}=\dim{\text{GBZ}}-\dim{E}$. Therefore, the dimension of the generalized Fermi surface may vary from 0 up to 2d22d-2. Unlike 1D non-Hermitian systems, which are characterized by a 0-dimensional generalized Fermi surface, systems with $d>1$ may exhibit $\dim{\text{GFS}}>0$. The presence of a nonzero $\dim{\text{GFS}}$ indicates that the summation in Eq. (17) becomes an integration under the thermodynamic limit with OBCs. Because integrals of exponential functions can engender nearly any function form (for instance, through the Laplace transform), even though the eigenstate is built from exponential functions of the coordinates $(x_{1},\ldots,x_{d})$ [Eq. (17)], it can manifest extremely diverse behaviors in non-Hermitian systems where $d>1$. For example, the eigenstate may turn into a power-law function of the coordinates, as demonstrated in Sec. IV. It is this richness that underpins the potential for the algebraic skin effect exclusive to non-Hermitian systems in dimensions higher than one.
@@
II.4 Strategy to formulate the universal GFS formula
Here, we succinctly outline our approach for deriving the GFS formula for non-Hermitian systems in arbitrary dimensions, with comprehensive formalism to be detailed in the subsequent sections. The GFS formula is used for constructing OBC wavefunctions in the thermodynamic limit, based on a standing-wave ansatz illustrated in Fig. 1. For finite-size lattice systems, we employ the layer transfer matrix method as a practical and efficient tool for computing the discretized GFS that matches the boundary degrees of freedom, with the standing-wave conditions automatically satisfied. In the thermodynamic limit, we derive continuous GFS curves analytically using the mathematical resultant method [32], as detailed in Appendix A, which agrees well with the finite-size results. To further demonstrate the applicability of the GFS framework, we apply it to a two-dimensional continuum model in Section VII, where an explicit analytic expression for the GFS is obtained without relying on the transfer matrix method. Below, we outline the layer transfer matrix approach for computing the GFS in finite-size lattice systems, while deferring the analytic treatment to Section VII and Appendix A.
Consider a system of dimension $d$, with the system size $L_{1}\times L_{2}\times\ldots\times L_{d}$. We can label the unit cells using integer coordinates $(x_{1},x_{2},\ldots,x_{3})$ with $1\leq x_{i}\leq L_{i}$. The transfer matrix method is employed to construct the eigenstates of the system — an exact analytic technique that systematically yields all eigenenergies and eigenstates for a given Hamiltonian. This method has been widely used for Hermitian systems [97, 98, 99] under both open and periodic boundary conditions, and it is equally effective for 1D non-Hermitian systems [100].
@@ II.4 Strategy to formulate the universal GFS formula
Here the bulk Hamiltonian is encoded in $\mathbb{T}$. For a fixed $\beta_{d}$, the reduced subsystem can be treated as a (d1)(d-1)-dimensional system, $\psi_{\beta_{d}}(x_{1},\ldots x_{d-1})$, and thus we can solve this (d1)(d-1)-dimensional problem using the transfer matrix method once again, defining a transfer matrix for $x_{d-1}$:
@@
Iterating this process ultimately simplifies the issue to a one-dimensional setup along the direction of $x_{1}$, $\psi_{\beta_{d}\beta_{d-1}\dots\beta_{2}}(x_{1})$. In this way, the $d$-dimensional system is reduced into a 1D system of $\psi_{\beta_{d}\beta_{d-1}\dots\beta_{2}}(x_{1})$, supplemented by transfer matrices for the other d1d-1 directions which only rely on bulk properties. We can then treat this 1D problem as a 1D non-Hermitian system, where the GBZ theory is well understood: the non-Hermitian skin effect necessitates a standing wave condition for the direction of $x_{1}$. This condition ensures that standing waves can be formed to obey boundary conditions at $x_{1}=1$ and $x_{1}=L_{1}$, allowing extensive number of skin modes. However, once this condition is enforced along $x_{1}$, imposing additional standing-wave constraints in other directions is generally untenable, which will be discussed below. Thus, for other directions, we must allow all possible eigenvalues of the transfer matrices $\mathbb{T}_{2},\ldots,\mathbb{T}_{d}$ without further stipulations.
@@
Since the standing wave condition can typically be imposed in just one spatial direction, as previously discussed, the GBZ’s complex wavevector space $(\beta_{1},\ldots,\beta_{d})$ characteristically spans a space of dimension 2d12d-1, surpassing real space dimensions $d$ if $d>1$. This distinctive characteristic underpins why non-Hermitian systems with dimensions greater than 1D can exhibit unprecedented phenomena.
@@
Figure 2:  Dimension analysis for GBZ and GFS in 1D and 2D systems. The GBZ dimension in $d$-dimensional systems ranges from $d$ to 2d12d-1. In 1D systems, $\dim{\text{GBZ}}=1$, and the OBC spectrum (denoted by $\sigma_{\text{OBC}}$) is limited to forming arcs, meaning $\dim{\sigma_{\text{OBC}}}=1$. This leads to $\dim{\text{GFS}}=0$ based on Eq. (5), with OBC eigenstates either being bulk or exponentially localized wavefunctions. In 2D systems, $\dim{\text{GBZ}}$ can be 22 or $3$, corresponding to the standing wave conditions that are enforced in two or one spatial dimensions, respectively. The OBC spectrum in 2D generally spans a finite area in the complex energy plane, $\dim{\sigma_{\text{OBC}}}=2$. Consequently, $\dim{\text{GBZ}}=0$ or $1$. When $\dim{\text{GFS}}=0$, the OBC eigenstates behave as bulk or exponentially localized wavefunctions. When $\dim{\text{GFS}}>0$, the OBC eigenstates can exhibit power-law decaying behavior, a phenomenon unique to high dimensions.
@@
As we conclude this section, it is important to draw attention to two key points. The first point is regarding the application of the standing wave condition, which can be imposed in an arbitrary direction. This flexibility implies that, depending on the choice of standing-wave-condition directions, $d$ distinct GBZs can be defined, and each GBZ corresponds to a unique set of basis functions. Because standing wave conditions in different directions in general produces different basis functions, they are typically not compatible with each other and thus cannot be simultaneously enforced. However, it is essential to recognize that while different bases associated with different GBZs can be used, the resulting eigenvalue spectrum and eigenstates—constructed as superpositions of these basis functions—remain invariant with respect to the choice of GBZ. Put simply, although multiple GBZ or GFS bases can be defined, they all ultimately describe the same physical reality. Secondly, although typically only one direction can accommodate the standing wave condition, special scenarios may allow the standing-wave conditions across multiple spatial directions, e.g., decoupled 1D chains populating a $d$-dimensional space. In such systems, GBZ dimension is further reduced to below 2d12d-1. If standing wave conditions for all directions are compatible with each other, the dimension of the GBZ reaches its minimum possible value, which coincides with the real-space dimensions $d$. This framework for one and higher dimensional systems is encapsulated in Fig. 1 and Fig. 2.
@@
III Generalized fermi surface formula in two dimensions based on the layer transfer matrix method
@@
In this section, we offer a general methodology for deriving the GFS formula of 2D lattice systems using a layer transfer matrix approach. The extension of the GFS formula to higher dimensions ($d>2$) is detailed in Section V; The application to a 2D continuum system is discussed in Section VII; The analytic approach to derive the GFS in the thermodynamic limit is provided in Appendix A.
@@
To construct OBC eigenstates, we begin with the layer transfer matrix for a given tight-binding Hamiltonian. The eigenvalues of the transfer matrix dictate the propagation of the wavefunction’s layer components along the transfer matrix direction. We demonstrate that a generic 2D OBC eigenstate can be represented by a superposition of standing wave components along one spatial direction (e.g., the $x$ direction), each coupled with the permissible non-Bloch wave $\beta_{y}$ in the $y$ direction. It is expressed as: $\psi(x,y)=\sum_{i}A_{i}\,\rho^{y}_{y,i}\varphi_{i}(x)$, where $\varphi_{i}(x)$ represents $i$-th standing-wave component along the $x$ direction and $\rho_{y,i}$ is the corresponding $i$-th transfer matrix eigenvalue. One of standing-wave components in a 2D OBC wavefunction is illustrated in Fig. 3. Here, two equal-amplitude, counter-propagating complex waves, $\beta_{x,i}$ and $\widetilde{\beta}_{x,i}$, form a standing wave $\varphi_{i}(x)$ along the $x$ direction and couple with the non-Bloch wave $\rho_{y,i}$ in the $y$ direction.
@@
Figure 3:  Schematic of $i$-th standing-wave component of an OBC eigenstate under the layer-$y$ transfer matrix basis. Here, the $i$-th component is a standing wave $\varphi_{i}(x)$ in the $x$ direction, formed by two equal amplitude and counter-propagating non-Bloch waves $\beta_{x,i}$ and $\widetilde{\beta}_{x,i}$, coupled with the associated $\rho_{y,i}$ in the $y$ direction.
As stated above, a set of basis functions to construct the OBC wavefunction can be determined by imposing the standing-wave condition along either the $x$ or $y$ spatial direction. This leads to distinct GFS or GBZ manifolds, depending on the chosen standing-wave direction. For example, in the cases of $n=1$ in two dimensions, as illustrated in Fig. 1, when applying standing-wave conditions in different directions, distinct GFS curves are obtained. A concrete example demonstrating this is provided in Appendix B. In contrast, when standing-wave conditions can be imposed along all directions, such as 1D systems and certain 2D cases, the GFS or GBZ is uniquely determined. Following this principle, the GFS formula applies to 2D systems with parallelogram-shaped open boundaries, where the entire lattice can be generated by translating a one-dimensional line of sites along a specific spatial direction.
In the following subsections, we elaborate on the main findings presented above and illustrate them through concrete examples. This section is organized as follows: Subsections III.1-III.3 focus on the GFS formalism for the most general cases in two dimensions, where standing waves can be imposed along only one direction and $\dim{\text{GBZ}}=3$, as illustrated by the 2D case with $n=1$ in Fig. 1. In Subsection III.4, we comment on the non-uniqueness of higher-dimensional GBZ and GFS manifolds and the associated standing-wave conditions. Finally, subsection III.5 discusses the other possible 2D scenarios.
III.1 Model and non-Bloch layer transfer matrix